Solve this system of equations and provide a graphical representation of the solution. (12)


x2 + y2 = 5


x + y = 1


1
Expert's answer
2022-05-24T11:51:57-0400

"\begin{cases}
 x^2+y^2=5 \\
 x+y=1
\end{cases}"

"\begin{cases}
 x^2+y^2=5 \\
 y=1-x
\end{cases}"

"x^2+(1-x)^2=5"

"x^2+1-2x+x^2=5"

"2x^2-2x-4=0"

"x^2-x-2=0"

"D=b^2-4ac=1+ 8=9"

"x_1=\frac{1+3}2=2"

"x_2=\frac{1-3}2=-1"

"y_1=-1"

"y_2=2"